OpenStudy (kitkat77):

5 • [(14 + 3 • 30) + 120] + 103 – 10 @Wendy.ivette11714

OpenStudy (wendy.ivette11714):

5(14+(3)(30)+120)+103−10 =5(14+90+120)+103−10 =5(104+120)+103−10 =(5)(224)+103−10 =1120+103−10 =1223−10 =1213

OpenStudy (kitkat77):

Thank you so much once more. You are a lifesaver.

OpenStudy (wendy.ivette11714):

Welcome